A particle of mass M1 moving with initial velocity Vo is incident on a stationary particle of mass M2. After collision , M1 was deflected through an angle ∅ and M2 an angle ∅. If the velocities of the particles after collision were V1 and V2 respectively. Show that for an elastic collission; V2=2Vo*M1/M1+M2*cos∅


1
Expert's answer
2021-02-18T07:17:49-0500
"m_1v_0=m_1v_1\cos{\theta}+m_2v_2\cos{\theta}\\m_1v_1\sin{\theta}=m_2v_2\sin{\theta}"

"m_1v_0=2m_2v_2\cos{\theta}\\v_2=\frac{m_1v_0}{2m_2\cos{\theta}}"